40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11181     Entry time: Sat Mar 28 03:21:49 2015
Author: den 
Type: Update 
Category: LSC 
Subject: towards angular ff 

Tonight I measured seismic noise coupling to beam spot on PR2. There is coherence of 0.9 from X to PIT and Y to YAW around the stack resonances. TF was fited using vectfit and put into static matrix of oaf in the elements T240X -> PRM PIT, T240Y -> PRM YAW. I think we should actuate on the error point of the PRM OL but I decided not to go for a model change tonight. Data from seismometers and POP QPD was obtained during the UTC time 04:06:00 - 04:50:00 when PRMI was locked on sideband

Interferometer was locking rather robustly and every lock lasted on the everage of 3 minutes. During these lock periods I incresed bandwidth of optical lever servos of BS and test masses from 4Hz up to 10Hz and then closed transmission QPD loops. It seems from the camera that lock losses correspond to strong motion of the beam on pop camera. Scripts that change OPLEV bandwidth are in /users/den "increase_ol_bandwidth.sh" "decrease_ol_bandwidth.sh". Script "engage_qpd_servos" turns off ETM oplevs and turns on ETM -> trans QPD servos. These scripts can be copied to locking directrly if are useful.

Please, note that transition from 3f to 1f should still be tuned. Only PRCL was stably controlled using 1f so far

ELOG V3.1.3-